OpenStudy (mikezack123):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

jimthompson5910 (jim_thompson5910):

point slope form y - y1 = m(x - x1) m is the slope (x1, y1) is the point the line goes through

OpenStudy (mikezack123):

isnt x and y where it goes through?

jimthompson5910 (jim_thompson5910):

no it's (x1,y1)

jimthompson5910 (jim_thompson5910):

x and y are left as they are

jimthompson5910 (jim_thompson5910):

the slope is 4, so m = 4

jimthompson5910 (jim_thompson5910):

the given point the line goes through is (-3, -1) so x1 = -3 y1 = -1

OpenStudy (mikezack123):

okay....i dont understand what to do with that information...

jimthompson5910 (jim_thompson5910):

you plug it into the equation y - y1 = m(x - x1)

jimthompson5910 (jim_thompson5910):

ie you replace m with 4, x1 with -3, y1 with -1 and you simplify

OpenStudy (mikezack123):

i got B

jimthompson5910 (jim_thompson5910):

good

OpenStudy (mikezack123):

WOOOOOOOOOOOOOOO!

OpenStudy (mikezack123):

thanks :)

jimthompson5910 (jim_thompson5910):

np